# Positive definiteness and Fell bundles over discrete groups

### v1

@@ 4 Some examples
Assume ${\mathcal{B}}=(B_{g})_{g\in G}$ is a Fell sub-bundle of a Fell Bundle ${\mathcal{A}}=(A_{g})_{g\in G}$, as defined in [27, Definition 21.5], and $E=(E_{g})$ is a conditional expectation from ${\mathcal{A}}$ to ${\mathcal{B}}$ in the sense of Exel [27, Definition 21.19]. Thus, each $E_{g}$ is a bounded idempotent linear mapping from $A_{g}$ onto $B_{g}$, and we have
@@
    $E_{gh}(ab)=E_{g}(a)\,b,\quad E_{hg}(ba)=bE_{g}(a).$
Then $E$ is a positive definite ${\mathcal{A}}$$-{\mathcal{B}}$ bundle map.
Indeed, if $g_{1},\ldots,g_{n}\in G$, $a_{i}\in A_{g_{i}}$ and $b_{i}\in B_{g_{i}}$ for $i=1,\ldots,n$, and we set $c:=\sum_{j=1}^{n}a_{j}b_{j}^{*}\in A_{e}$, then
$\sum_{i,j=1}^{n}b_{i}\,E_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\,b_{j}^{*}$ $=\sum_{i,j=1}^{n}\,E_{g_{i}g_{i}^{-1}g_{j}g_{j}^{-1}}(b_{i}a_{i}^{*}a_{j}b_{j}^{*})$
$=\sum_{i,j=1}^{n}\,E_{e}(b_{i}a_{i}^{*}a_{j}b_{j}^{*})=E_{e}(c^{*}c)$
@@
for all $g\in G$ and $a\in A_{g}$. It is well-known that such a map $M_{E}$ exists and is a conditional expectation when $C_{r}^{*}({\mathcal{B}})$ is identified with its canonical copy inside $C_{r}^{*}({\mathcal{A}})$ (cf. [27, Theorem 21.29]).
Further, since ${\mathcal{B}}$ admits a conditional expectation in the sense of Exel (by assumption), we can identify $C^{*}({\mathcal{B}})$ with its canonical copy inside $C^{*}({\mathcal{A}})$ (cf. [27, Theorem 21.30]). Then, for every $g\in G$ and $b\in B_{g}$, we get that
$\Phi_{E}(\widehat{j}^{\,{\mathcal{A}}}_{g}(b))=\widehat{j}^{\,{\mathcal{A}}}_{g}(E_{g}(b))=\widehat{j}^{\,{\mathcal{A}}}_{g}(b).$
Hence, by continuity of $\Phi_{E}$, we get that $\Phi_{E}(x)=x$ for every $x\in C^{*}({\mathcal{B}})$, i.e., $\Phi_{E}$ is a projection from $C^{*}({\mathcal{A}})$ onto $C^{*}({\mathcal{B}})$. Since $\|\Phi_{E}\|=\|E_{e}\|=1$, we get that $\Phi_{E}$ is a conditional expectation.
@@
Let $Y$ be a right Hilbert $C_{r}^{*}({\mathcal{B}})$-module and $y\in Y$. Assume that $C_{r}^{*}({\mathcal{A}})$ acts on $Y$ from the left by adjointable operators. For each $g\in G$, define a linear map $T_{g}:A_{g}\to B_{\varphi(g)}$ by
$T_{g}(a)=E^{\mathcal{B}}_{\varphi(g)}\Big{(}\big{\langle}y,\lambda^{\mathcal{A}}_{g}(a)y\big{\rangle}_{C_{r}^{*}({\mathcal{B}})}\Big{)}\quad\text{for all }a\in A_{g}.$ (5)
Then $T=(T_{g})$ is clearly an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. Moreover, it is positive definite:
Indeed, let $g_{1},\ldots,g_{n}\in G$ and $a_{i}\in A_{g_{i}},b_{i}\in B_{\varphi(g_{i})}$ for $i=1,\ldots,n$. Then we have
$b_{i}\,T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\,b_{j}^{*}$ $=b_{i}\,E^{\mathcal{B}}_{\varphi(g_{i}^{-1}g_{j})}\big{(}\langle y,\lambda^{\mathcal{A}}_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})y\rangle_{C_{r}^{*}({\mathcal{B}})}\big{)}\,b_{j}^{*}$
$=b_{i}\,E^{\mathcal{B}}_{\varphi(g_{i})^{-1}\varphi(g_{j})}\big{(}\langle\lambda^{\mathcal{A}}_{g_{i}}(a_{i})y,\lambda^{\mathcal{A}}_{g_{j}}(a_{j})y\rangle_{C_{r}^{*}({\mathcal{B}})}\big{)}b_{j}^{*}$
$=b_{i}\,E^{\mathcal{B}}_{\varphi(g_{i})^{-1}\varphi(g_{j})}\big{(}z_{ij}\big{)}b_{j}^{*},$
where $z_{ij}:=\langle\lambda^{\mathcal{A}}_{g_{i}}(a_{i})y,\lambda^{\mathcal{A}}_{g_{j}}(a_{j})y\rangle_{C_{r}^{*}({\mathcal{B}})}$ for each $i,j$. Now, using [27, Proposition 17.12], we get that
$b_{i}\,E^{\mathcal{B}}_{\varphi(g_{i})^{-1}\varphi(g_{j})}\big{(}z_{ij}\big{)}b_{j}^{*}=\lambda^{\mathcal{B}}_{g_{i}}(b_{i})\,z_{ij}\,\lambda^{\mathcal{B}}_{g_{j}}(b_{j})^{*}$
for each $i,j$. Since $Z:=[z_{ij}]\in M_{n}(C_{r}^{*}({\mathcal{B}}))^{+}$, we get that
$\sum_{i,j=1}^{n}b_{i}\,T_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\,b_{j}^{*}=\sum_{i,j=1}^{n}\lambda^{\mathcal{B}}_{g_{i}}(b_{i})\,z_{ij}\,\lambda^{\mathcal{B}}_{g_{j}}(b_{j})^{*}=\big{\langle}\Omega,Z\,\Omega\big{\rangle}\in C_{r}^{*}({\mathcal{B}})^{+}\,,$
[…]
We leave to the reader to check that one obtains a similar result if one replaces the reduced cross-sectional $C^{*}$-algebras by their full counterparts and defines $T_{g}:A_{g}\to B_{\varphi(g)}$ by
$T_{g}(a)=\big{(}E^{\mathcal{B}}_{\varphi(g)}\circ\Lambda^{\mathcal{B}}\big{)}\Big{(}\big{\langle}y,\widehat{j}^{\mathcal{A}}_{g}(a)y\big{\rangle}_{C^{*}({\mathcal{B}})}\Big{)}\quad\text{for all }a\in A_{g}.$ (6)
Conversely, let us start with a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$. We can then form the completely positive map $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$. Following Paschke [55, Section 5] (see also [46, p. 48]), we may then construct from $\Phi_{T}$ a right Hilbert $C^{*}({\mathcal{B}})$-module $Y$ and a left action of $C^{*}({\mathcal{A}})$ on $Y$ by adjointable operators. Assume that $A_{e}$ and $B_{e}$ are unital. Then $C^{*}({\mathcal{A}})$ and $C^{*}({\mathcal{B}})$ are unital, so [55, Theorem 5.2] gives that there exists some $y\in Y$ such that
@@
for all $X\in C^{*}({\mathcal{A}})$. Then for all $g\in G$ and $a\in A_{g}$ we have
@@
This shows that every positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T$ may be written in the form (6). If we further assume that ${\rm ker}(\varphi)$ is amenable, then arguing analogously with the completely positive map $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$, we also get that every such bundle map $T$ is of the form (5).
@@ 5 An approximation property for Fell bundles
Let $\Sigma=(A,G,\alpha)$ be a (unital) discrete $C^{*}$-dynamical system, and let ${\mathcal{B}}$ denote the associated Fell bundle over $G$. If the system $\Sigma$ is amenable in the sense of [10], then, making use of Example 4.7, we get that ${\mathcal{B}}$ has the PD-approximation property.
@@
The proof of the first assertion is an adaptation of the proof for the case of group $C^{*}$-algebras associated to amenable groups given in [14, Theorem 2.6.8] (see also the proof of [10, Theorem 4.6]).
We start by observing that if $F$ is a finite subset of $G$ and $C_{F}$ is the subspace of $C_{c}({\mathcal{B}})$ given by $C_{F}=\{f\in C_{c}({\mathcal{B}}):{\rm supp}(f)\subseteq F\}$, then we have that $\kappa(C_{F})$ is a closed subspace of $C^{*}({\mathcal{B}})$. Here we write $\kappa$ for the canonical map $\kappa^{\mathcal{B}}:C_{c}({\mathcal{B}})\to C^{*}({\mathcal{B}})$.
Indeed, let $\{f_{n}\}$ be a sequence in $C_{F}$ such that $\{\kappa(f_{n})\}$ converges to some $x\in C^{*}({\mathcal{B}})$. For each $g\in G$, let $\widetilde{E}_{g}:C^{*}({\mathcal{B}})\to B_{g}$ be the linear map defined by $\widetilde{E}_{g}:=E^{\mathcal{B}}_{g}\circ\Lambda^{\mathcal{B}}$, which is clearly continuous. For each $g\in F$, we then have
$f_{n}(g)=\widetilde{E}_{g}(\kappa(f_{n}))\to\widetilde{E}_{g}(x)\in B_{g}\,\,\text{as}\,\,n\to\infty\,.$
Thus, letting $f\in C_{F}$ be defined by $f=\sum_{g\in F}\widetilde{E}_{g}(x)\odot g$, and using (1), we get
$\|x-\kappa(f)\|_{u}$ $\leq\|x-\kappa(f_{n})\|_{u}+\|\kappa(f_{n}-f)\|_{u}$
$\leq\|x-\kappa(f_{n})\|_{u}+\sum_{g\in F}\|f_{n}(g)-\widetilde{E}_{g}(x)\|\to 0\,\,\text{as}\,\,n\to\infty\,.$
Hence, $x=\kappa(f)\in\kappa(C_{F})$, as desired.
Next, let $x\in C^{*}({\mathcal{B}})$. Consider a positive definite ${\mathcal{B}}$-bundle map $T$ having finite support. Then $\Phi_{T}(x)\in\kappa(C_{c}({\mathcal{B}}))$: indeed, letting $F$ denote the support of $T$ and $\{f_{n}\}$ be a sequence in $C_{c}({\mathcal{B}})$ such that $\{\kappa(f_{n})\}$ converges to $x$, we have
@@
Assume now that $\Lambda(x)=0$. Let $\{T^{i}\}_{i\in I}$ be a net as guaranteed by the PD-approximation property of ${\mathcal{B}}$. Theorem 3.10 (resp. Theorem 3.12) gives a net $\{\Phi_{T^{i}}\}_{i\in I}$ (resp. $\{M_{T^{i}}\}_{i\in I}$) of completely positive maps on $C^{*}({\mathcal{B}})$ (resp.  $C_{r}^{*}({\mathcal{B}})$). Note that for each $i\in I$, we have $\Lambda\circ\Phi_{T^{i}}=M_{T^{i}}\circ\Lambda$ on $C^{*}({\mathcal{B}})$, since the maps on both sides are continuous and agree on $\kappa(C_{c}({\mathcal{B}}))$. Hence, for each $i\in I$, we get
@@
Since each $T^{i}$ is finitely supported, we have $\Phi_{T^{i}}(x)\in\kappa(C_{c}({\mathcal{B}}))$, as established above. But $\Lambda$ is injective on $\kappa(C_{c}({\mathcal{B}}))$, cf. [27, Proposition 17.9], so we obtain that $\Phi_{T^{i}}(x)=0$ for each $i\in I$.
@@
Hence, using the assumption that $\lim_{i}\|T^{i}_{g}(b)-b\|=0$ for every $g\in G$ and $b\in B_{g}$, we get that $\lim_{i}\|\Phi_{T^{i}}(\kappa(f))-\kappa(f)\|_{\rm u}=0$. This implies that $\lim_{i}\|\Phi_{T^{i}}(y)-y\|_{\rm u}=0$ for all $y\in\kappa\big{(}C_{c}({\mathcal{B}})\big{)}$. If $\{u_{j}\}_{j\in J}$ is an approximate unit for $A$, then $\{\kappa\big{(}u_{j}\odot e)\}_{j\in J}$ is easily seen to be an approximate unit for $C^{*}({\mathcal{B}})$, so we get that
@@
Our next result involves the tensor product Fell bundle $C\otimes_{\rm min}{\mathcal{B}}=\big{(}C\otimes_{\rm min}B_{g}\big{)}_{g\in G}$, which is defined for any $C^{*}$-algebra $C$ in [27, Definition 25.4]. A similar result when ${\mathcal{B}}$ has the AP is shown by Exel in [27, Proposition 25.9].
@@
We recall from [27, Sections 24 and 25] that the Fell bundle $\mathcal{D}:=C\otimes_{\rm min}{\mathcal{B}}$ is the completion of the pre-Fell-bundle obtained by equipping the algebraic Fell bundle $C\odot{\mathcal{B}}=(C\odot B_{g})_{g\in G}$ with the minimal $C^{*}$-norm $\|\cdot\|_{\rm min}$ on $C\odot B_{e}$. The norm on each fiber $C\odot B_{g}$ is then given by $\|d\|=\|d^{*}d\|_{\rm min}^{1/2}$ for $d\in C\odot B_{g}$, and the fiber of $C\otimes_{\rm min}{\mathcal{B}}$ at $g$, denoted by $C\otimes_{\rm min}B_{g}$, is the completion of $C\odot B_{g}$ w.r.t. this norm.
@@
for all $c\in C,g\in G$ and $b\in B_{g}$, and $\|\widetilde{M}_{i}\|=\|{\rm id}_{C}\|~\|M_{T^{i}}\|=\|T^{i}_{e}\|$. Now, using [27, Theorem 25.8], there is a *-isomorphism $\psi:C_{r}^{*}(\mathcal{D})\to C\otimes_{\rm min}C_{r}^{*}({\mathcal{B}})$ satisfying
@@
By Theorem 5.4, we know that $C^{*}({\mathcal{B}})\simeq C_{\rm r}^{*}({\mathcal{B}})$. Assume that $B_{e}$ is nuclear. In [27, Proposition 25.10], Exel shows that $C^{*}({\mathcal{B}})\simeq C_{\rm r}^{*}({\mathcal{B}})$ is then nuclear whenever ${\mathcal{B}}$ has the AP. It is straightforward to see that his proof goes through verbatim if one just invokes Proposition 5.5 instead of [27, Proposition 25.9]. Conversely, nuclearity of $B_{e}$ is necessary for the nuclearity of $C_{r}^{*}({\mathcal{B}})$ since there exists a conditional expectation from $C_{r}^{*}({\mathcal{B}})$ onto $B_{e}$. $\square$
We can now include the PD-approximation property in [3, Proposition 7.2] (see also [5, Théorème 4.5], [27, Proposition 25.10] and [16, Corollary 4.23]).

### v2

@@ 4 Some examples
Assume ${\mathcal{B}}=(B_{g})_{g\in G}$ is a Fell sub-bundle of a Fell Bundle ${\mathcal{A}}=(A_{g})_{g\in G}$, as defined in [28, Definition 21.5], and $E=(E_{g})_{g\in G}$ is a conditional expectation from ${\mathcal{A}}$ to ${\mathcal{B}}$ in the sense of Exel [28, Definition 21.19]. Thus, each $E_{g}$ is a bounded idempotent linear mapping from $A_{g}$ onto $B_{g}$, and we have
@@
    $E_{g}(a)^{*}=E_{g^{-1}}(a^{*}),\quad E_{gh}(ab)=E_{g}(a)\,b,\quad E_{hg}(ba)=bE_{g}(a).$
Then $E$ is a positive definite ${\mathcal{A}}$$-{\mathcal{B}}$ bundle map. Indeed, if $g_{1},\ldots,g_{n}\in G$, $a_{i}\in A_{g_{i}}$ and $b_{i}\in B_{g_{i}}$ for $i=1,\ldots,n$, and we set $c:=\sum_{j=1}^{n}a_{j}b_{j}^{*}\in A_{e}$, then
$\sum_{i,j=1}^{n}b_{i}\,E_{g_{i}^{-1}g_{j}}(a_{i}^{*}a_{j})\,b_{j}^{*}=\sum_{i,j=1}^{n}\,E_{g_{i}g_{i}^{-1}g_{j}g_{j}^{-1}}(b_{i}a_{i}^{*}a_{j}b_{j}^{*})=\sum_{i,j=1}^{n}\,E_{e}(b_{i}a_{i}^{*}a_{j}b_{j}^{*})=E_{e}(c^{*}c)$
@@
for all $g\in G$ and $a\in A_{g}$. It is well-known that such a map $M_{E}$ exists and is a conditional expectation when $C_{r}^{*}({\mathcal{B}})$ is identified with its canonical copy inside $C_{r}^{*}({\mathcal{A}})$ (cf. [28, Theorem 21.29]). In the present situation, we may also identify $C^{*}({\mathcal{B}})$ with its canonical copy inside $C^{*}({\mathcal{A}})$ (cf. [28, Theorem 21.30]), and it is not difficult to check that $\Phi_{E}$ is a conditional expectation as well.
@@
Let $Y$ be a right Hilbert $C_{r}^{*}({\mathcal{B}})$-module and $y\in Y$. Assume that $C_{r}^{*}({\mathcal{A}})$ acts on $Y$ from the left by adjointable operators. As is well-known and easy to check, we then get a completely positive linear map $M:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ by setting
$M(X)=\big{\langle}y,X\cdot y\big{\rangle}_{C_{r}^{*}({\mathcal{B}})}\quad\text{for all }X\in C_{r}^{*}({\mathcal{A}}).$
By Proposition 3.8, we get a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ by setting
$T_{g}(a):=E^{\mathcal{B}}_{\varphi(g)}\Big{(}M\big{(}\lambda^{\mathcal{A}}_{g}(a)\big{)}\Big{)}=E^{\mathcal{B}}_{\varphi(g)}\Big{(}\big{\langle}y,\lambda^{\mathcal{A}}_{g}(a)\cdot y\big{\rangle}_{C_{r}^{*}({\mathcal{B}})}\Big{)}\quad\text{for all }a\in A_{g}.$ (4)
Replacing the reduced cross-sectional $C^{*}$-algebras by their full counterparts and using Proposition 3.9, we get a similar result by defining $T_{g}:A_{g}\to B_{\varphi(g)}$ by
$T_{g}(a)=\big{(}E^{\mathcal{B}}_{\varphi(g)}\circ\Lambda^{\mathcal{B}}\big{)}\Big{(}\big{\langle}y,\widehat{j}^{\mathcal{A}}_{g}(a)\cdot y\big{\rangle}_{C^{*}({\mathcal{B}})}\Big{)}\quad\text{for all }a\in A_{g}.$ (5)
Conversely, let us start with a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$. We can then form the completely positive map $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$. Following Paschke [55, Section 5] (see also [47, p. 48]), we may use $\Phi_{T}$ to construct a right Hilbert $C^{*}({\mathcal{B}})$-module $Y$ and a left action of $C^{*}({\mathcal{A}})$ on $Y$ by adjointable operators. Assume that $A_{e}$ and $B_{e}$ are unital. Then $C^{*}({\mathcal{A}})$ and $C^{*}({\mathcal{B}})$ are unital, so [55, Theorem 5.2] gives that there exists some $y\in Y$ such that
@@
for all $X\in C^{*}({\mathcal{A}})$. Then for all $g\in G$ and $a\in A_{g}$, we have
@@
This shows that when ${\mathcal{A}}$ and ${\mathcal{B}}$ are unital, every positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T$ may be written in the form (5). If we further assume that ${\rm ker}(\varphi)$ is amenable, then arguing analogously with $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$, we also get that every such bundle map $T$ is of the form (4).
@@ 5 An approximation property for Fell bundles
Let $\Sigma=(A,G,\alpha)$ be a unital discrete $C^{*}$-dynamical system, and let ${\mathcal{B}}$ denote the associated Fell bundle over $G$. If the system $\Sigma$ is amenable in the sense of [10], then, making use of Example 4.6, we get that ${\mathcal{B}}$ has the PD-approximation property.
@@
The proof is an adaptation of the proof for the case of group $C^{*}$-algebras associated to amenable groups given in [14, Theorem 2.6.8] (see also the proof of [10, Theorem 4.6]).
We start by observing that if $F$ is a finite subset of $G$ and $C_{F}$ is the subspace of $C_{c}({\mathcal{B}})$ given by $C_{F}=\{f\in C_{c}({\mathcal{B}}):{\rm supp}(f)\subseteq F\}$, then it is not difficult to check that $\kappa(C_{F})$ is a closed subspace of $C^{*}({\mathcal{B}})$. Here we write $\kappa$ for the canonical map $\kappa^{\mathcal{B}}:C_{c}({\mathcal{B}})\to C^{*}({\mathcal{B}})$.
Next, let $x\in C^{*}({\mathcal{B}})$. Consider a positive definite ${\mathcal{B}}$-bundle map $T$ having finite support. Then $\Phi_{T}(x)\in\kappa(C_{c}({\mathcal{B}}))$: indeed, letting $F$ denote the support of $T$ and $\{f_{n}\}$ be a sequence in $C_{c}({\mathcal{B}})$ such that $\{\kappa(f_{n})\}$ converges to $x$, we have (using the notation introduced in Section 2)
@@
Assume now that $\Lambda(x)=0$, where $\Lambda:=\Lambda^{\mathcal{B}}$ is the canonical *-homomorphism from $C^{*}({\mathcal{B}})$ onto $C^{*}_{r}({\mathcal{B}})$. Let $\{T^{i}\}_{i\in I}$ be a net as guaranteed by the PD-approximation property of ${\mathcal{B}}$. Theorem 3.11 (resp. Theorem 3.14) gives a net $\{\Phi_{T^{i}}\}_{i\in I}$ (resp. $\{M_{T^{i}}\}_{i\in I}$) of completely positive maps on $C^{*}({\mathcal{B}})$ (resp.  $C_{r}^{*}({\mathcal{B}})$). Note that for each $i\in I$, we have $\Lambda\circ\Phi_{T^{i}}=M_{T^{i}}\circ\Lambda$ on $C^{*}({\mathcal{B}})$, since the maps on both sides are continuous and agree on $\kappa(C_{c}({\mathcal{B}}))$. Hence, for each $i\in I$, we get
@@
Since each $T^{i}$ is finitely supported, we have $\Phi_{T^{i}}(x)\in\kappa(C_{c}({\mathcal{B}}))$, as established above. But $\Lambda$ is injective on $\kappa(C_{c}({\mathcal{B}}))$, cf. [28, Proposition 17.9], so we obtain that $\Phi_{T^{i}}(x)=0$ for each $i\in I$.
@@
Hence, using the assumption that $\lim_{i}\|T^{i}_{g}(b)-b\|=0$ for every $g\in G$ and $b\in B_{g}$, we get that $\lim_{i}\|\Phi_{T^{i}}(\kappa(f))-\kappa(f)\|_{\rm u}=0$. This implies that $\lim_{i}\|\Phi_{T^{i}}(y)-y\|_{\rm u}=0$ for all $y\in\kappa\big{(}C_{c}({\mathcal{B}})\big{)}$. If $\{u_{j}\}_{j\in J}$ is an approximate unit for $B_{e}$, then $\{\kappa\big{(}u_{j}\odot e)\}_{j\in J}$ is easily seen to be an approximate unit for $C^{*}({\mathcal{B}})$, so we get that
@@
Our next result involves the tensor product Fell bundle $C\otimes_{\rm min}{\mathcal{B}}=\big{(}C\otimes_{\rm min}B_{g}\big{)}_{g\in G}$, which is defined for any $C^{*}$-algebra $C$ in [28, Definition 25.4]. A similar result when ${\mathcal{B}}$ has the AP is shown by Exel in [28, Proposition 25.9].
@@
We recall from [28, Sections 24 and 25] that the Fell bundle $\mathcal{D}:=C\otimes_{\rm min}{\mathcal{B}}$ is the completion of the pre-Fell-bundle obtained by equipping the algebraic Fell bundle $C\odot{\mathcal{B}}=(C\odot B_{g})_{g\in G}$ with the minimal $C^{*}$-norm $\|\cdot\|_{\rm min}$ on $C\odot B_{e}$. The norm on each fiber $C\odot B_{g}$ is then given by $\|d\|=\|d^{*}d\|_{\rm min}^{1/2}$ for $d\in C\odot B_{g}$, and the fiber of $C\otimes_{\rm min}{\mathcal{B}}$ at $g$, denoted by $C\otimes_{\rm min}B_{g}$, is the completion of $C\odot B_{g}$ w.r.t. this norm.
@@
for all $c\in C,g\in G$ and $b\in B_{g}$, and $\|\widetilde{M}_{i}\|=\|{\rm id}_{C}\|~\|M_{T^{i}}\|=\|T^{i}_{e}\|$. Now, using [28, Theorem 25.8], there is a *-isomorphism $\psi:C_{r}^{*}(\mathcal{D})\to C\otimes_{\rm min}C_{r}^{*}({\mathcal{B}})$ satisfying
@@
By Theorem 5.4, we know that $C^{*}({\mathcal{B}})\simeq C_{\rm r}^{*}({\mathcal{B}})$. Assume that $B_{e}$ is nuclear. In [28, Proposition 25.10], Exel shows that $C^{*}({\mathcal{B}})\simeq C_{\rm r}^{*}({\mathcal{B}})$ is then nuclear whenever ${\mathcal{B}}$ has the AP. It is straightforward to see that his proof goes through verbatim if one just invokes Proposition 5.5 instead of [28, Proposition 25.9]. Conversely, nuclearity of $B_{e}$ is necessary for the nuclearity of $C_{r}^{*}({\mathcal{B}})$ since there exists a conditional expectation from $C_{r}^{*}({\mathcal{B}})$ onto $B_{e}$. $\square$
We can now include the PD-approximation property in [3, Proposition 7.2] (see also [5, Théorème 4.5], [28, Proposition 25.10] and [16, Corollary 4.23]).